# SKALD: Scalable K-Anonymisation for Large Datasets

### v1

@@ I Introduction
In the modern era, data is a pseudo-currency that underpins digital and social transactions. However, the data economy is fraught with significant privacy risks [Dwork2017][Abowd2023] that threaten individual confidentiality. This makes safeguarding the privacy of data containing personal and sensitive information a critical concern if the data is to be shared, analysed or released.
Globally, regulatory pressures drive the data anonymisation imperative, and advisory guidelines are provided on techniques and execution of anonymisation [Gadotti2024]. In India, the Digital Personal Data Protection (DPDP) Act [GovernmentofIndia2023] is a precursor to more stringent and clearly outlined legislation around anonymisation. Canada, the EU and the UK stipulate that data should be rendered anonymous such that the data subject is no longer identifiable, while Singapore delineates acceptable privacy parameter ranges for the sharing of anonymised data through well-established privacy-preserving techniques such as k-anonymisation [Sweeney2002][DSCI2023].
The General Data Protection Regulation (GDPR), long considered the gold standard for comprehensive data protection rules, outlines stringent guidelines that emphasise the importance of data sovereignty [EuropeanParliament2016a]. The transfer of personal data to third parties or international organisations is strictly regulated. Local processing of data is encouraged to ensure an auditable chain of control and accountability which implies the need for dedicated hardware systems to carry out anonymisation tasks. Available hardware resources provide an upper bound on the maximum size of dataset that can be processed at a time. Large datasets with sizes exceeding this upper bound must be broken up into smaller data chunks for processing. Although optimal k-anonymisation implementation requires knowledge of the entire dataset, standard k-anonymisation tools such as ARX can only operate on a per-chunk basis if the dataset is too large. This lack of a global view of the dataset can negatively affect the output information utility.
@@ III-A Bounding the Number of Histogram Bins
The amount of available physical memory (RAM) constrains the amount of space available for a histogram of a dataset. Equivalently, this also imposes a bound on the maximum number of bins in the histogram, which we denote by $N_{RAM}$. Each bin in our histogram represents a specific combination of multi-dimensional values across the QIDs.
Consider a dataset $D$ with $L$ categorical and $M$ numerical QIDs. For the $i^{th}$ categorical QID with resolution $R_{i}$, let $C_{i}[R_{i}]$ denote the number of unique categorical values. For numerical QIDs, we assume the existence of metadata that provides us with the maximum and minimum values. For the $j^{th}$ numerical QID, let $Q_{j,max}$, $Q_{j,min}$ and $W_{j}$ denote the maximum value, minimum value, and bin width respectively. The number of bins in the histogram are given by
$N_{data}=\Big{[}(\prod_{i=1}^{L}C_{i}[R_{i}])(\prod_{j=1}^{M}(Q_{j,max}-Q_{j,min}+1)/W_{j})\Big{]}+1$ (1)
where,
  • $\prod_{i=1}^{L}C_{i}[R_{i}]$ denotes the number of unique combinations of values across categorical QIDs in the dataset for a given resolution
  • $\prod_{j=1}^{M}(Q_{j,max}-Q_{j,min}+1)/W_{j}$ denotes the number of unique combinations of values across numerical QIDs with respectively chosen bin widths
  • The addition of the one to the product of both the above terms represents the suppressed records, which constitute their own bin in the histogram
The upper bound on the number of bins of a histogram due to the RAM constraint is given by
$N_{data}\leq N_{RAM}$ (2)
The resolution $R_{i}$ for categorical QIDs and the bin widths $W_{j}$ for numerical QIDs should be chosen to retain as much fine-grained information as possible, while satisfying (2). The exact procedure for selecting these parameters is described below.
@@ III-B2 Phase 2: Histogram and Lattice Construction
In the second phase, the data is passed chunk-wise through the algorithm for a second time. A histogram is created to store the counts of users that fall into the bins defined by $(R_{i,\mathrm{RAM}},W_{j,\mathrm{RAM}})$. This histogram is updated for each chunk, so the final histogram created after passing all the chunks has the aggregate counts from all the chunks, i.e. the entire dataset. A new generalisation lattice is then constructed, using $(R_{i,\mathrm{RAM}},W_{j,\mathrm{RAM}})$ as the root node. For all nodes that pass the k-anonymity condition within the desired suppression limit, we compute the monotonic Discernibility Metric (DM*) proposed in [ElEmam2009]. This enables us to select the node with the lowest (best) DM* value, corresponding to minimal information loss, among all the nodes that satisfy k-anonymity. We label this node $(R_{i,\mathrm{final}},W_{j,\mathrm{final}})$.
@@ V Results
A representative snippet of the output generated by SKALD for $k$ = 50 using 125 chunks is shown in Table II. This configuration resulted in the $(R_{i,\mathrm{final}},W_{j,\mathrm{final}})$ pair: ((2, 1), (1, 24, 32)).
TABLE III: Results for $k=1000$ across varying no. of chunks
@@
1 $((3,1),(4,24,1347))$ $((3,3),(67,16,4))$
5 $((3,1),(1,24,1347))$ $((3,3),(67,16,4))$
25 $((3,1),(1,24,256))$ $((3,3),(67,16,4))$
125 $((3,1),(1,24,128))$ $((3,3),(67,16,4))$
Table III compares the generalisation nodes of SKALD and ARX for different numbers of chunks when $k=1000$. SKALD algorithm extracts and combines sufficient statistics from each chunk to provide reduced bin widths with an increase in the number of chunks, where each data chunk contains 1 million records. This reduced bin width retains more fine-grained information. ARX, on the other hand, anonymises the data on a per-chunk basis and shows no improvement in the generalisation node with an increasing number of chunks.

### v2

@@ I Introduction
In the modern era, data is a pseudo-currency that forms the foundation of digital and social transactions. However, the data economy is fraught with significant privacy risks [Dwork2017][Abowd2023] that threaten individual confidentiality. This makes safeguarding the privacy of data containing private and sensitive data a key issue if the data is to be shared, analysed or released.
Globally, regulatory pressures drive the data anonymisation imperative, and advisory guidelines are provided on techniques and execution of anonymisation [Gadotti2024]. In India, the Digital Personal Data Protection (DPDP) Act [GovernmentofIndia2023] is a precursor to more stringent and clearly outlined legal frameworks around anonymisation. Canada, the EU and the UK stipulate that data should be rendered anonymous such that the data subject is no longer identifiable, while Singapore delineates acceptable privacy parameter ranges for the distribution of anonymised data through well-established privacy-preserving techniques such as k-anonymisation [Sweeney2002][DSCI2023].
The General Data Protection Regulation (GDPR), long considered the gold standard for extensive data security measures, outlines stringent guidelines that emphasise the significance of data autonomy [EuropeanParliament2016a]. The transfer of personal data to third parties or international organisations is strictly regulated. Local processing of data is encouraged to ensure an auditable chain of oversight and responsibility which implies the need for dedicated hardware systems to carry out anonymisation tasks. The available hardware resources provide an upper bound on the maximum size of dataset that can be processed at a time. Large datasets with sizes exceeding this upper bound must be broken up into smaller data chunks for processing. Although optimal k-anonymisation implementation requires knowledge of the entire dataset, standard k-anonymisation tools such as ARX can only operate on a per-chunk basis if the dataset is too large. This lack of a global view of the dataset can negatively affect the utility of the output information.
@@ III-A Bounding the Number of Histogram Bins
The amount of available physical memory (RAM) constrains the space available for a histogram of a dataset. Equivalently, this imposes a bound on the maximum number of bins in the histogram, denoted by $N_{RAM}$. Each bin corresponds to one equivalence class—a unique combination of values across all QIDs.
Consider a dataset $D$ with $L$ categorical and $M$ numerical QIDs. For the $i$th categorical QID at resolution $R_{i}$, let $C_{i}[R_{i}]$ be the number of unique categorical values. For each numerical QID $j$, let $Q_{j,\max}$, $Q_{j,\min}$, and $W_{j}$ denote its maximum value, minimum value, and chosen bin width, respectively. For these parameters, the total number of bins (equivalence classes) in the histogram, denoted by $N_{data}$, is given by
[…]
where:
  • $\prod_{i=1}^{L}C_{i}[R_{i}]$ is the number of unique combinations of categorical QID values at resolutions $\{R_{i}\}$.
  • $\prod_{j=1}^{M}\bigl{(}(Q_{j,\max}-Q_{j,\min}+1)/W_{j}\bigr{)}$ is the number of numerical‐QID bins at widths $\{W_{j}\}$.
  • The “$+1$” term accounts for suppressed records as a separate bin.
The parameters $\{R_{i}\}$ and $\{W_{j}\}$ in (1) are chosen to satisfy
$N_{data}\;\leq\;N_{RAM}.$ (2)
Next, let $n$ be the number of records per chunk and $d$ be the size (in bytes) of a single record. Then the chunk occupies
$S\;=\;n\times d\quad(\text{bytes of RAM}).$
As SKALD operates by extracting and combining sufficient statistics across all chunks, the accuracy of SKALD does not depend on the chunk size. We can ensure that the storage of histograms in SKALD does not require additional memory by operating with smaller chunks. In particular, we can use a reduced chunk size of $n/2$ records occupying $S/2$ bytes, and the remaining $S/2$ bytes can be split between two histograms—i.e. each histogram can use at most $S/4$ bytes. Since each bin in a histogram requires a 4 byte integer to store the count, the maximum number of bins per histogram satisfies
$N_{RAM}\;=\;\frac{S/4}{4}\;=\;\frac{S}{16}.$
Substituting $S=n\times d$ yields the RAM‐based upper bound
$N_{RAM}\;=\;\frac{n\times d}{16}.$ (3)
Equations (2) and (3) together force
$N_{data}\;\leq\;\frac{n\times d}{16}.$
Therefore, when choosing categorical resolutions $\{R_{i}\}$ and numerical bin widths $\{W_{j}\}$, one must retain as much granularity as possible while ensuring
[…]
The detailed procedure for selecting $\{R_{i}\}$ and $\{W_{j}\}$ under this constraint is described in the subsequent sections.
@@ III-B2 Phase 2: Histogram and Lattice Construction
In the second phase, the data is passed chunk-wise through the algorithm for a second time. A histogram is created to store the counts of users that fall into the bins defined by $(R_{i,\mathrm{RAM}},W_{j,\mathrm{RAM}})$. This histogram is updated for each chunk, so the final histogram created after processing all the chunks contains aggregate counts over the entire dataset.
A new generalisation lattice is then constructed, using $(R_{i,\mathrm{RAM}},W_{j,\mathrm{RAM}})$ as the root node. Note that we have the knowledge of the histogram for the root node, i.e. $(R_{i,\mathrm{RAM}},W_{j,\mathrm{RAM}})$, containing aggregate counts over the entire dataset. To apply predictive tagging, we proceed by checking the $k$-anonymity condition from the middle node of the middle level in the generalisation lattice. This requires merging some bins in the root histogram in order to construct the histogram corresponding to the chosen node. We maintain an unmodified copy of the root histogram, and as per the node’s generalisation level we merge the bins in the root histogram to create a new histogram. We now evaluate whether the new histogram satisfies the $k$-anonymity condition. After evaluating a node and marking it as “pass” or “fail”, we proceed to evaluate the next node in the generalisation lattice.
For all nodes that pass the $k$-anonymity condition within the desired suppression limit, we compute the monotonic Discernibility Metric (DM*) proposed in [ElEmam2009]. This enables us to select the node with the lowest (best) DM* value, corresponding to minimal information loss, among all the nodes that satisfy $k$-anonymity. We label this node $(R_{i,\mathrm{final}},W_{j,\mathrm{final}})$.
@@ V Results
A representative snippet of the output generated by SKALD for $k$ = 50 using 125 chunks is shown in Table II. This configuration resulted in the generalization node [$(R_{i,\mathrm{final}},W_{j,\mathrm{final}})$]: ((2, 1), (1, 24, 32)).
TABLE III: Results for $k=50$ across varying no. of chunks
@@
5 $((3,1),(1,24,128))$ $((3,4),(4,24,4))$
25 $((3,1),(1,24,32))$ $((3,4),(4,24,4))$
125 $((2,1),(1,24,32))$ $((3,4),(4,24,4))$
Table III compares the generalisation nodes of SKALD and ARX for different numbers of chunks when $k=50$. SKALD algorithm extracts and combines sufficient statistics from each chunk to provide reduced bin widths with an increase in the number of chunks, where each data chunk contains 1 million records. This reduced bin width retains more fine-grained information. ARX, on the other hand, anonymises the data on a per-chunk basis and shows no improvement in the generalisation node with an increasing number of chunks.